# Global Convergence of Policy Gradient for Entropy Regularized Linear-Quadratic Control with Multiplicative Noises

### v3

@@ Abstract
Reinforcement Learning (RL) has emerged as a powerful framework for sequential decision-making in dynamic environments, particularly when system parameters are unknown. This paper investigates RL-based control for entropy-regularized linear-quadratic (LQ) control problems with multiplicative noises over an infinite time horizon. First, we adapt the regularized policy gradient (RPG) algorithm to stochastic optimal control settings, proving that despite the non-convexity of the problem, RPG converges globally under conditions of gradient domination and near-smoothness. Second, based on zero-order optimization approach, we introduce a novel model free RL algorithm: Sample-based regularized policy gradient (SB-RPG). SB-RPG operates without knowledge of system parameters yet still retains strong theoretical guarantees of global convergence. Our model leverages entropy regularization to address the exploration versus exploitation trade-off inherent in RL. Numerical simulations validate the theoretical results and demonstrate the efficacy of SB-RPG in unknown-parameters environments.
@@ 1 Introduction
In the context of entropy-regularized RL for LQ problems, most research has focused on actor-critic methods [23, 24, 18], with relatively few studies addressing policy gradient approaches. Notably, Michael et al. [7] study the global linear convergence of policy gradient methods for finite-horizon continuous-time entropy-regularized LQ control problems. In discrete time settings, Guo et, al. [10] proposed and analyzed two new policy gradient based RL method for entropy-regularized LQ problems: regularized policy gradient (RPG) and iterative policy optimization (IPO) and proved their fast convergence given exact model parameters(i.e., model based). However, they only considered additive noises and conducted their analysis solely in the model-based setting, which limits the applicability of their methods. Multiplicative noise models may be produce more robust policies, they are still more complex than the typical additive model which may result in slower convergence. To the best of our knowledge, addressing the case of multiplicative noise in both model-based and model-free settings still remains largely unexplored.
This paper makes two primary contributions: First, we extend the RPG [10] method to stochastic LQ scenarios with multiplicative noises, demonstrating that RPG achieves global convergence despite the non-convexity of the problem, thanks to gradient domination and almost-smoothness properties. This extension enhances robustness and broadens the applicability of RPG. Second, and more importantly, employing the zero order optimization technique, we propose sample based regularized policy gradient (SB-RPG) method and rigorously prove its global convergence properties. SB-RPG does not require knowledge of the specific system parameters values, enabling effective control in parameter-unknown environments where RPG cannot be directly applied. Our theoretical results are supported by numerical simulations.
@@ 2 Formulation
where $A,C,x_{t},w_{t}^{x}\in\mathbb{R},~B\in\mathbb{R}^{1\times n},u,w_{t}^{u}\in\mathbb{R}^{n\times 1},~D\in\mathbb{R}^{n\times n},$ and $w_{t}^{x},w_{t}^{u}$ are white noises, which are distributed as follows
@@ 4 Global Convergence of Sample Based Regularized Policy Gradient
$l\geq\frac{\log\epsilon-\log S_{K,\Sigma}}{log\gamma}$
@@
$l\geq\frac{\log\epsilon-\log\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right]}{log\gamma}$
@@
$\|\nabla_{K}f(K^{\prime},\Sigma^{\prime})-\nabla_{K}f(K,\Sigma)\|\leq h_{6}\|K^{\prime}-K\|+h_{7}\|\Sigma^{\prime}-\Sigma\|$
@@
$\|\nabla_{\Sigma}f(K^{\prime},\Sigma^{\prime})-\nabla_{\Sigma}f(K,\Sigma)\|\leq h_{8}\|K^{\prime}-K\|+h_{9}\|\Sigma^{\prime}-\Sigma\|$
@@
We define $f_{r_{1}}(K,\Sigma):=\mathbb{E}_{U\sim\mathbb{B}_{r_{1}}}[f(K+U,\Sigma)],$ and $f_{r_{2}}(K,L):=\mathbb{E}_{V\sim\mathbb{B}_{r_{2}}}[f(K,(L+V)(L+V)^{\top}]$. where $\mathbb{B}_{r}$ denotes the uniform distribution over the points with norm $r$ (boundary of a sphere). The following lemma shows that the gradient of $f_{r_{1}}(K,\Sigma)$ and $f_{r_{2}}(K,L)$ can be estimated with an oracle for the function value.
@@
$\nabla_{K}f_{r_{1}}(K,\Sigma)=\frac{n}{r_{1}^{2}}\mathbb{E}_{U\sim\mathbb{S}_{r_{1}}}[f(K+U,\Sigma)U].$
@@
$\nabla_{L}f_{r_{2}}(K,L)=\frac{n(n+1)}{2r_{2}^{2}}\mathbb{E}_{V\sim\mathbb{S}_{r_{2}}}[f(K,(L+V)(L+V)^{\top})V].$
@@
Assume that (i) the distribution of the initial states implies that $\|x_{0}^{i}\|\leq\bar{L}$ almost surely for any $i$. (ii) the multiplicative noises are distributed such that $\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{l-1}Qx_{t}^{2}+u_{t}^{\top}Ru_{t}+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\right]$ for any $i$. If set
$r_{1}$ $\leq\frac{\epsilon_{1}}{2h_{6}}$
$M$ $\geq\max\big\{\frac{2n}{(\epsilon_{1}/6)^{2}}(\sigma_{1}+\frac{R_{1}\epsilon}{18\sqrt{n}})\log\big(\frac{n+1}{\sqrt{\kappa_{1}}}\big),\frac{2n}{(\epsilon/3)^{2}}(\sigma^{2}_{2}+\frac{R_{2}\epsilon}{9x\sqrt{n}})\log(\frac{n+1}{\sqrt{\kappa_{1}}})\big\}$
$l$ $\geq\log(\gamma)^{-1}\left[\log\left(\frac{r_{1}}{n}\cdot\frac{\epsilon}{3}\right)-\log\left(2|f(K,\Sigma)|\left(2\|K\|^{2}\|R\|+\frac{1}{|Q|}\right)+|\psi|\left(1+\frac{1}{|Q|}+\frac{1}{1-\gamma}\right)\right)\right]$
@@
Assume that (a) the distribution of the initial states implies that $\|x_{0}^{i}\|\leq\bar{L}$ almost surely for any $i$. (b) the multiplicative noises are distributed such that $\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}|x_{t})\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\right].$ If
$r_{2}$ $\leq\frac{\epsilon_{2}}{12h_{9}\|L^{-1}\|}$
$M$ $\geq\frac{2n(n+1)}{(\epsilon_{2}/3)^{2}}(\sigma_{Z}^{2}+\frac{R_{Z}{\epsilon_{2}/3}}{3})\log\big(\frac{n(n+1)/2+1}{1-\sqrt{1-\kappa}}\big)$
$l$ $\geq\frac{\log\frac{2r_{2}\epsilon_{2}}{6n(n+1)\|L^{-1}\|}-\log\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right]}{\log\gamma}$
@@
$|f(K,\Sigma)-f(K^{\prime},\Sigma^{\prime})|\leq h_{10}\|K^{\prime}-K\|+h_{11}\|\Sigma^{\prime}-\Sigma\|$
@@
Given an arbitrary tolerance $\epsilon>0$ and probability $\kappa\in(0,1)$. If we set
@@ Appendix A Proofs in Section 2
$=(2RK+\gamma P_{K}\nabla V_{K})\sum_{t=0}^{\infty}(\gamma V_{K})^{t}.$
@@ Appendix B Proofs in Section 3
$+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$
@@
with equality when $K^{\prime}-K=-(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}$
@@
Let $S_{K}=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}[x_{t}^{2}]$ and remember that $\nabla_{K}f(K,\Sigma)=S_{K}E_{K}$
@@
$\leq\frac{S_{K^{*},\Sigma^{*}}}{\mu^{2}\sigma_{min}(R)}\nabla^{\top}_{K}f_{K,\Sigma}(x_{0})\nabla_{K}f_{K,\Sigma}(x_{0})+\frac{(1-\gamma)\text{Tr}[(\nabla_{\Sigma}C_{K,\Sigma}(x_{0}))^{2}]}{\sigma_{min}(R)}$
@@
$f(K,\Sigma)-f(K^{*},\Sigma)\leq\frac{1}{\mu\sigma_{min}(R)}\nabla_{K}f^{\top}(K,\Sigma)\nabla_{K}f(K,\Sigma)+\frac{(1-\gamma)\|\nabla_{\Sigma}f(K,\Sigma)\|^{2}}{\sigma_{min}(R)}$ (14)
@@
taking the expectation w.r.t. $x_{0}$ on both sides we have,
@@
$\leq\frac{f(K,\Sigma)-\Omega}{Q}\sqrt{\lambda_{1}^{-1}(f(K,\Sigma)-f(K^{*},\Sigma^{*}))}$ (15)
@@
$\leq(1-\gamma)^{-1}\left(\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|+\frac{\tau}{2\sigma_{min}(\Sigma)}\right)$
@@
Observe:
$q_{K,\Sigma}-q_{K,\Sigma^{\prime}}+Tr\left(\nabla_{\Sigma}q_{K,\Sigma}^{\top}(\Sigma^{\prime}-\Sigma)\right)=\frac{\tau}{2(1-\gamma)}\left[log(\Sigma^{-1}\Sigma^{\prime})-Tr(\Sigma^{-1}\Sigma^{\prime}-I)\right]$ (17)
@@
We will first show:
@@
$\leq-\frac{\eta_{2}}{2(1-\gamma)^{2}}\text{Tr}[(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{2}]$
$\eta_{2}\leq\frac{2(1-\gamma)a^{2}}{\tau}\leq\frac{2(1-\gamma)}{\tau}\left(\frac{\tau}{2\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}\right)\leq\frac{2(1-\gamma)}{\tau}\leq\frac{(1-\gamma)}{\tau m}$
@@
$\leq\frac{1}{(1-\gamma)a\sigma_{min}(R)}Tr[((R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}I)^{2}]$
@@
$q_{K,\Sigma^{\prime}}-q_{K,\Sigma}\leq\frac{\eta_{2}a\sigma_{min}(R)}{2(1-\gamma)}(q_{K,\Sigma}-q_{K,\Sigma^{*}})$
@@
$f(K^{\prime},\Sigma^{\prime})-f(K^{*},\Sigma^{*})\leq(1-\phi)(f(K,\Sigma)-f(K^{*},\Sigma^{*}))$
@@
$q_{K,\Sigma}\geq\frac{\tau k}{2(1-\gamma)}log\left(\frac{\sigma_{min}(R)}{\pi\tau}\right)$

### v4

@@ Abstract
Reinforcement Learning (RL) has emerged as a powerful framework for sequential decision-making in dynamic environments, particularly when system parameters are unknown. This paper investigates RL-based control for entropy-regularized linear-quadratic (LQ) control problems with multiplicative noise over an infinite time horizon. First, we adapt the regularized policy gradient (RPG) algorithm to stochastic optimal control settings, proving that despite the non-convexity of the problem, RPG converges globally under conditions of gradient domination and almost-smoothness. Second, based on zero-order optimization approach, we introduce a novel model free RL algorithm: Sample-based regularized policy gradient (SB-RPG). SB-RPG operates without knowledge of system parameters yet still retains strong theoretical guarantees of global convergence. Our model leverages entropy regularization to address the exploration versus exploitation trade-off inherent in RL. Numerical simulations validate the theoretical results and demonstrate the efficiency of SB-RPG in unknown-parameters environments.
@@ 1 Introduction
In the context of entropy-regularized RL for LQ problems, most research has focused on actor-critic methods [23, 24, 18], with relatively few studies addressing policy gradient approaches. Notably, Michael et al. [7] study the global linear convergence of policy gradient methods for finite-horizon continuous-time entropy-regularized LQ control problems. In discrete time settings, Guo et, al. [10] proposed and analyzed two new policy gradient based RL method for entropy-regularized LQ problems: regularized policy gradient (RPG) and iterative policy optimization (IPO) and proved their fast convergence given exact model parameters(i.e., model based). However, they only considered additive noise and conducted their analysis solely in the model-based setting, which limits the applicability of their methods. Multiplicative noise models may be produce more robust policies, they are still more complex than the typical additive model which may result in slower convergence. To the best of our knowledge, addressing the case of multiplicative noise in both model-based and model-free settings still remains largely unexplored.
This paper makes two primary contributions: First, we extend the RPG [10] method to stochastic LQ scenarios with multiplicative noise, demonstrating that RPG achieves global convergence despite the non-convexity of the problem, thanks to gradient domination and almost-smoothness properties. This extension enhances robustness and broadens the applicability of RPG. Second, and more importantly, employing the zero order optimization technique, we propose sample based regularized policy gradient (SB-RPG) method and rigorously prove its global convergence properties. SB-RPG does not require knowledge of the specific system parameters values, enabling effective control in parameter-unknown environments where RPG cannot be directly applied. Our theoretical results are supported by numerical simulations.
@@ 2 Formulation
where $A,C,x_{t},w_{t}^{x}\in\mathbb{R},~B\in\mathbb{R}^{1\times n},u,w_{t}^{u}\in\mathbb{R}^{n\times 1},~D\in\mathbb{R}^{n\times n},$ and $w_{t}^{x},w_{t}^{u}$ are white noise, which are distributed as follows
@@ 4 Global Convergence of Sample Based Regularized Policy Gradient
$l\geq\frac{\log\epsilon-\log S_{K,\Sigma}}{log\gamma}.$
@@
$l\geq\frac{\log\epsilon-\log\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right]}{log\gamma}.$
@@
$\|\nabla_{K}f(K^{\prime},\Sigma^{\prime})-\nabla_{K}f(K,\Sigma)\|\leq h_{6}\|K^{\prime}-K\|+h_{7}\|\Sigma^{\prime}-\Sigma\|,$
@@
$\|\nabla_{\Sigma}f(K^{\prime},\Sigma^{\prime})-\nabla_{\Sigma}f(K,\Sigma)\|\leq h_{8}\|K^{\prime}-K\|+h_{9}\|\Sigma^{\prime}-\Sigma\|,$
@@
We define $f_{r_{1}}(K,\Sigma):=\mathbb{E}_{U\sim\mathbb{B}_{r_{1}}}[f(K+U,\Sigma)]$ and $f_{r_{2}}(K,L):=\mathbb{E}_{V\sim\mathbb{B}_{r_{2}}}[f(K,(L+V)(L+V)^{\top}]$. where $\mathbb{B}_{r}$ denotes the uniform distribution over the points with norm $r$ (boundary of a sphere). The following lemma shows that the gradient of $f_{r_{1}}(K,\Sigma)$ and $f_{r_{2}}(K,L)$ can be estimated with an oracle for the function value.
@@
$\nabla_{K}f_{r_{1}}(K,\Sigma)=\frac{n}{r_{1}^{2}}\mathbb{E}_{U\sim\mathbb{S}_{r_{1}}}[f(K+U,\Sigma)U],$
@@
$\nabla_{L}f_{r_{2}}(K,L)=\frac{n(n+1)}{2r_{2}^{2}}\mathbb{E}_{V\sim\mathbb{S}_{r_{2}}}[f(K,(L+V)(L+V)^{\top})V],$
@@
Assume that (i) the distribution of the initial states implies that $\|x_{0}^{i}\|\leq\bar{L}$ almost surely for any $i$. (ii) the multiplicative noise are distributed such that $\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{l-1}Qx_{t}^{2}+u_{t}^{\top}Ru_{t}+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\right]$ for any $i$. If set
$r_{1}$ $\leq\frac{\epsilon_{1}}{2h_{6}},$
$M$ $\geq\max\big\{\frac{2n}{(\epsilon_{1}/6)^{2}}(\sigma_{1}+\frac{R_{1}\epsilon}{18\sqrt{n}})\log\big(\frac{n+1}{\sqrt{\kappa_{1}}}\big),\frac{2n}{(\epsilon/3)^{2}}(\sigma^{2}_{2}+\frac{R_{2}\epsilon}{9x\sqrt{n}})\log(\frac{n+1}{\sqrt{\kappa_{1}}})\big\},$
$l$ $\geq\log(\gamma)^{-1}\left[\log\left(\frac{r_{1}}{n}\cdot\frac{\epsilon}{3}\right)-\log\left(2|f(K,\Sigma)|\left(2\|K\|^{2}\|R\|+\frac{1}{|Q|}\right)+|\psi|\left(1+\frac{1}{|Q|}+\frac{1}{1-\gamma}\right)\right)\right],$
@@
Assume that (a) the distribution of the initial states implies that $\|x_{0}^{i}\|\leq\bar{L}$ almost surely for any $i$. (b) the multiplicative noise are distributed such that $\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}|x_{t})\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\right].$ If
$r_{2}$ $\leq\frac{\epsilon_{2}}{12h_{9}\|L^{-1}\|},$
$M$ $\geq\frac{2n(n+1)}{(\epsilon_{2}/3)^{2}}(\sigma_{Z}^{2}+\frac{R_{Z}{\epsilon_{2}/3}}{3})\log\big(\frac{n(n+1)/2+1}{1-\sqrt{1-\kappa}}\big),$
$l$ $\geq\frac{\log\frac{2r_{2}\epsilon_{2}}{6n(n+1)\|L^{-1}\|}-\log\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right]}{\log\gamma},$
@@
$|f(K,\Sigma)-f(K^{\prime},\Sigma^{\prime})|\leq h_{10}\|K^{\prime}-K\|+h_{11}\|\Sigma^{\prime}-\Sigma\|,$
@@
Given an arbitrary tolerance $\epsilon>0$ and probability $\kappa\in(0,1)$. If
@@ Appendix A Proofs in Section 2
$=(2RK+\gamma P_{K}\nabla V_{K})\sum_{t=0}^{\infty}(\gamma V_{K})^{t},$
@@ Appendix B Proofs in Section 3
$+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}}).$
@@
with equality when $K^{\prime}-K=-(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}$.
@@
Let $S_{K}=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}[x_{t}^{2}]$ and remember that $\nabla_{K}f(K,\Sigma)=S_{K}E_{K}$.
@@
$\leq\frac{S_{K^{*},\Sigma^{*}}}{\mu^{2}\sigma_{min}(R)}\nabla^{\top}_{K}f_{K,\Sigma}(x_{0})\nabla_{K}f_{K,\Sigma}(x_{0})+\frac{(1-\gamma)\text{Tr}[(\nabla_{\Sigma}C_{K,\Sigma}(x_{0}))^{2}]}{\sigma_{min}(R)}.$
@@
$f(K,\Sigma)-f(K^{*},\Sigma)\leq\frac{1}{\mu\sigma_{min}(R)}\nabla_{K}f^{\top}(K,\Sigma)\nabla_{K}f(K,\Sigma)+\frac{(1-\gamma)\|\nabla_{\Sigma}f(K,\Sigma)\|^{2}}{\sigma_{min}(R)}.$ (14)
@@
. taking the expectation w.r.t. $x_{0}$ on both sides we have
@@
$\leq\frac{f(K,\Sigma)-\Omega}{Q}\sqrt{\lambda_{1}^{-1}(f(K,\Sigma)-f(K^{*},\Sigma^{*}))},$ (15)
@@
$\leq(1-\gamma)^{-1}\left(\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|+\frac{\tau}{2\sigma_{min}(\Sigma)}\right).$
@@
Observe
$q_{K,\Sigma}-q_{K,\Sigma^{\prime}}+Tr\left(\nabla_{\Sigma}q_{K,\Sigma}^{\top}(\Sigma^{\prime}-\Sigma)\right)=\frac{\tau}{2(1-\gamma)}\left[log(\Sigma^{-1}\Sigma^{\prime})-Tr(\Sigma^{-1}\Sigma^{\prime}-I)\right]$
@@
We will first show that
@@
$\leq-\frac{\eta_{2}}{2(1-\gamma)^{2}}\text{Tr}[(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{2}],$
$\eta_{2}\leq\frac{2(1-\gamma)a^{2}}{\tau}\leq\frac{2(1-\gamma)}{\tau}\left(\frac{\tau}{2\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}\right)\leq\frac{2(1-\gamma)}{\tau}\leq\frac{(1-\gamma)}{\tau m},$
@@
$\leq\frac{1}{(1-\gamma)a\sigma_{min}(R)}Tr[((R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}I)^{2}].$
@@
$q_{K,\Sigma^{\prime}}-q_{K,\Sigma}\leq\frac{\eta_{2}a\sigma_{min}(R)}{2(1-\gamma)}(q_{K,\Sigma}-q_{K,\Sigma^{*}}).$
@@
and
@@
$q_{K,\Sigma}\geq\frac{\tau k}{2(1-\gamma)}log\left(\frac{\sigma_{min}(R)}{\pi\tau}\right).$